OpenStudy (anonymous):

the sum of three times a number and -4 is at least twice the number plus 8 define a variable, write an inequality, and solve each problem

OpenStudy (kamibug):

We can use the variable "a." Here is the inequality: 3a - 4 >= 2a + 8 Here are the steps to solve it: 3a - 4 >= 2a + 8 3a >= 2a + 12 a >= 12

OpenStudy (anonymous):

thank you!

OpenStudy (kamibug):

Sure :)